Ch.22 #11

A 1.62-m-long FM antenna is oriented parallel to the electric field of an EM wave.

Part A

How large must the electric field be to produce a 1.05-mV (rms) voltage between the ends of the antenna?

Part B

What is the rate of energy transport per square meter?
Answer #1

Solution)

Part a)

We know,

Electric Field, E=V(rms)/d

= 1.05*10^-3/1.62 m

=0.000648 V/m

========

Part b)

Now,

Rate of Energy=?

We know, Electruc Field, E =c*εo*E^2

Here,

E = 0.000648 V/m

So,

Rate of energy= 3*10^8*8.854*10^-12*(0.000648)^2

= 1.15*10^-9 W/m^2

===============